A long solenoid (cross-sectional area = 1.10*10^-6 m^2, number of turns per unit length = 2544 turns/m) is bent into a circular shape so it looks like a donut. This wire-wound donut is called a toroid. Assume that the diameter of the solenoid is small compared to the radius of the toroid, which is 0.052 m. Find the emf induced in the toroid when the current decreases to 1.3 A from 2.7 A in a time of 0.14 s.

The answer is 2.92*10^-5. Just need to know the steps to solve. Thanks!
Answer #1

e = -L*dI/dt

we know

L = uo*n^2*A*length

length = circumference of circle = 2pi*R

n = 2544 turns/m

A = 1.10 x 10^-6

R = 0.052 m

dI = (1.3 - 2.7) = -1.4

dt = 0.14

uo = 4pi x 10^-7

e = 2.92 x 10^-7 V